(3x+4) by (1x+2)
1
Expert's answer
2020-07-29T12:01:56-0400

(3x+4)/(1x+2)

3x+4=1x+2

3x-1x=2-4

2x=-2

x=-1

Put these values in the given equation

(3x+4)/(1x+2)

((3(-1)+4))/(-1+2)

(-3+4)/(-1+2)

1